OpenStudy (anonymous):

How many solutions does this system have? 2x + y = 3 6x = 9 -3y please explain

OpenStudy (anonymous):

None.

OpenStudy (akshay_budhkar):

GT explain man :P

OpenStudy (dumbcow):

2x + y = 3 6x +3y = 9 multiply top equation by -3 to cancel out x's -6x -3y = -9 6x +3y = 9 ------------ 0 = 0 Everything cancels, since 0=0 is True there are infinite solutions This is because the 2 equations are essentially the same line so any point (x,y) on the line is a solution.